Penny is making punch for Sam's party. The recipe requires twice as much orange juice as cranberry juice and 8 times as much ginger ale as cranberry juice. if she plans to make 176 ounces of punch, how much of each ingredient should she use?
Answer by ptaylor(2198)

Let x=amount of cranberry juice needed
Then 2x=amount of orange juice needed
And 8x=amount of ginger ale needed
Now our equation to solve is:
x+2x+8x=176 collect like terms
11x=176 divide each side by 11
x=16 oz-----------------amount of cranberry juice needed
2x=2*16=32 oz------------amount of orange juice needed
8x=8*16= 128 oz-------------amount of ginger ale needed
Ck
16+32+128=176
176=176


Hope this helps---ptaylor